# MSGNav: Unleashing the Power of Multi-modal 3D Scene Graph for Zero-Shot Embodied Navigation

### v1

@@ 1 Introduction
Previous RL-based embodied navigation methods [rlnav1, rlnav2] suffer from poor generalization and a large sim-to-real gap [sg-nav]. Inspired by rapid advances in pre-trained large model technique [gpt4o, qwen], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [3dmem, dynavlm, vlngame, tango] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
@@
  • Visual Deficiency. Converting visual observations into text-only graphs discards visual evidence, leading to non-robustness against perception errors.
  • Constrained Vocabulary: Novel categories beyond a preset vocabulary cannot be represented, limiting generalization in 3D scene graph-based methods.
Figure 2: Performance comparison between our approaches with others. (a). The superiority of our M3DSG over traditional 3D scene graphs. (b). Distance statistics from the goal for the previous method (3D-Mem [3dmem] as example). (c). Our MSGNav navigation system achieves state-of-the-art performance on both the HM3D-OVON and Goat-Bench benchmarks.
Based on these analyses, we conclude that visual information is indispensable to the navigation process. Therefore, we introduce the Multi-modal 3D Scene Graph (M3DSG) to address these limitations. As illustrated in Fig. 1, M3DSG replaces the pure-text relational edges with dynamically assigned images. Our M3DSG effectively addresses the three limitations above:
@@
  • Visual Replenishment: M3DSG provides visual context to enhance robustness against perception errors.
  • Unconstrained Vocabulary: M3DSG utilizes preserved visual context to expand vocabulary, enabling open-vocabulary generalization dynamically.
@@
Additionally, we also realize that the last‑mile problem in navigation tasks — knowing the target’s location does not equate to determining a suitable navigation pose [^1] Cannot see the wood for the trees. Being too close or occluded prevents a comprehensive observation of target, thus is not a good viewpoint., and prior embodied navigation methods stop before that final step. As shown in Fig. 2(b), the previous method [3dmem] successfully realizes the target but still fails many navigation tasks due to unacceptable viewpoints (distance ranging from 0.25 m to 1.0 m). To address this, MSGNav includes a visibility‑based viewpoint decision module.
@@
  • M3DSG: We propose a multi-modal 3D scene graph that integrates visual information, overcoming pure-text limitations and enhancing open-vocabulary scene representation for embodied navigation.
  • Last-Mile: We identify the last-mile problem—selecting optimal viewpoints for the final target approach. Meanwhile, we propose a visibility‑based viewpoint decision module to address it.
  • MSGNav: We develop a zero-shot embodied navigation system built on M3DSG, enabling active exploration and achieving state-of-the-art performance on GOAT-Bench and HM3D-OVON.
@@ 3 Approach
Figure 3: The overall framework of our MSGNav. At time step $t$, the agent incrementally constructs the scene graph $S_{t}$ based on received observation $\mathcal{I}_{t}$ and its own pose. $S_{t}$ includes a set of objects $\textbf{O}_{t}$ with attributes such as visual, spatial, and room properties, along with a set of image edges $\textbf{E}_{t}$ representing relationships between objects. Subsequently, $S_{t}$ is processed through KSS, AVU, and CLR modules, before being input to VLM for querying to obtain the target object $\bar{o}$. Finally, VVD module selects the insightful viewpoint $\textbf{v}_{best}$ as a navigation point.
@@ 3.1.1 Problem definition
We consider a mobile embodied agent operating in an unknown or partially explored environment(lifelong navigation setting). The agent receives target information in the form of an category, natural language description, or reference image. At each time step $t$, it obtains an RGB-D observation $\mathcal{I}_{t}$ and executes an action $\mathcal{A}_{t}$ (camera rotation or ego-motion) to actively explore until locating the target. The task is successful if the agent reaches any target viewpoint within $d$ meters in at most $T$ steps; otherwise, it fails. In the zero-shot navigation setting, the agent performs navigation without task-specific training or fine-tuning in simulation, thereby reducing training cost and improving generalization to unseen environments and targets.
@@ 3.1.2 Overview
Zero-shot goal-oriented embodied navigation remains highly challenging for conventional learning-based methods. Following current best practices, we utilize off-the-shelf pre-trained models: Vision Foundation Models (VFMs) and Vision Language Models (VLMs) for primitive reasoning. While these models excel in visual understanding, they are limited in 3D spatial reasoning. To address this gap, we propose M3DSG, an explicit, incremental, and multimodal 3D scene graph that encodes scene context while preserving visual evidence. Building on M3DSG, the navigation system MSGNav completes goal-oriented navigation with a Visibility-based Viewpoint Decision (VVD) module.
@@ 3.2 M3DSG: Multi‑modal 3D Scene Graph
Traditional 3D scene graphs represent object relations with text-only edges, which are overly abstract and hinder VLMs from accurately understanding scene context. Inspired by 3D-Mem [3dmem], which emphasizes the value of raw images for navigation, we retain visual information when constructing an object-centric 3D scene graph. Unlike traditional 3D scene graph  [Conceptgraphs], which uses textual relation edges, our method stores images to describe detailed object relations directly. This image edge preserving the benefits of 3D scene graphs while avoiding repeated, costly model queries and providing a more holistic scene representation.
@@ 3.2.1 M3DSG Structure Definition
During the process of exploring the environment, M3DSG constructs a scene graph $\mathbf{S}=(\mathbf{O},\mathbf{E})$, where $\mathbf{O}=\{o_{i}\}_{i=1}^{N_{o}}$ and $\mathbf{E}=\{\mathbf{e}_{j}\}_{j=1}^{N_{e}}$ denote the sets of objects and edges, respectively. Here, $N_{o}$ is the number of detected objects, and $N_{e}=\sum_{1\leq x<y\leq N_{o}}\mathds{1}_{\|Pos_{o_{x}}-Pos_{o_{y}}\|<\theta}$ is the number of object pairs $(o_{x},o_{y})$ whose spatial distance is below $\theta$. Each object $o_{i}$ is described by attributes such as category, 3D coordinates, and room location. Each edge $\mathbf{e}_{j}$ stores a set of RGB-D images $\mathbf{I}_{j}=\{\mathcal{I}_{k}\}_{k=1}^{|\mathbf{I}_{j}|}$ that represent the pair of corresponding objects to record their relationship.
@@ 3.2.2 Incremental Construction of M3DSG
Following prior work [sg-nav, unigoal], the scene graph $\mathbf{S}$ is incrementally updated during exploration. At step $t$, $\mathbf{S}_{t-1}$ is updated to $\mathbf{S}_{t}$ with the current RGB-D observation $\mathcal{I}_{t}$:
@@
Object Update. For the RGB-D observation $\mathcal{I}_{t}$ at time $t$, we extract the frame-level object set $\mathbf{O}^{frame}_{t}=\{o_{i}\}_{i=1}^{|\mathbf{O}^{frame}_{t}|}$ by VFMs: YOLO-W [yolo] for open-vocabulary detection, SAM [sam] for instance masks, and CLIP [clip] for visual embeddings. Each object $o_{i}$ is represented as $o_{i}=\{\mathcal{ID},\mathcal{C},\mathcal{P},\mathcal{B},\mathcal{M},\mathcal{PC},\mathcal{V},\mathcal{R}\}_{i},$ denoting unique ID, category, 3D position, bounding box, semantic mask, point cloud, visual feature, and room location, respectively. The global object set $\mathbf{O}_{t}$ is incrementally updated from previous global object set $\mathbf{O}_{t-1}$ and frame-level object set $\mathbf{O}^{frame}_{t}$:
@@
Edge Update. The core of an edge is a set of images $\mathbf{I}$ in which an object pair co-occurs, indexed by the unique IDs of the object pair: $\mathbf{E}:(\mathcal{ID},\mathcal{ID})\rightarrow\mathbf{I}.$
Meanwhile, we also maintain a mapping from each image to its corresponding object pair: $\mathbf{H}:\mathcal{I}\rightarrow(\mathcal{ID},\mathcal{ID}).$
For the current RGB-D frame $\mathcal{I}_{t}$ at step $t$ with object set $\mathbf{O}^{frame}_{t}$ (from Object Update), we generate co-occurring object pairs $(o_{x},o_{y})$ that appear together in $\mathcal{I}_{t}$:
@@
where $\theta$ is the adjacency distance threshold. The edge set $\mathbf{E}_{t}$ is then updated by retrieving the unique IDs $(\mathcal{ID}_{o_{x}},\mathcal{ID}_{o_{y}})$ of each detected pair and appending $\mathcal{I}_{t}$ to the corresponding image set:
@@
This edge update process is efficient, eliminating the need for costly VLM queries. Moreover, the stored image records in $\mathbf{E}_{t}$ provide compact, visually grounded representations of object relationships, which can be directly leveraged for downstream exploration reasoning.
@@ 3.3 MSGNav Embodied Navigation System
M3DSG provides scene representations that preserve rich visual information. To fully exploit this, we propose the embodied navigation system MSGNav (Multi-modal 3D Scene Graph for Navigation). As shown in Fig.  3(b), MSGNav first significantly reduces the cost of tokens and time required for inference by selecting key subgraphs (Sec. 3.3.1). MSGNav then addresses the out-of-vocabulary issue inherent to pre-set vocabulary by adaptively updating vocabulary with visual evidence from M3DSG (Sec. 3.3.2). Meanwhile, MSGNav performs closed-loop reasoning via decision memory and feedback reasoning (Sec. 3.3.3). Finally, MSGNav tackles the “last-mile” problem by optimizing navigation points through visibility-based viewpoint decisions (Sec. 3.3.4).
@@ 3.3.1 Key Subgraph Selection (KSS)
During exploration, the 3D scene graph $\mathbf{S}$ grows progressively and can become voluminous, thereby impairing decision-making efficiency [3dmem]. In embodied navigation, however, only a small related subgraph to the required target is typically meaningful. Therefore, processing the full scene graph consumes substantial computational resources without commensurate benefit. To address this, the objective is to preserve maximal relevant information with the minimal scene graph. MGSNav extracts the target‑relevant subgraphs via a Compress–Focus–Prune procedure:
@@
7:for $o\in\mathbf{O}$ do $\triangleright$ Filter relevant edges
8:  if $\exists~{o}^{r}\in\textbf{O}^{rel}$: $(\mathcal{ID}_{o},\,\mathcal{ID}_{{o}^{r}})\in\mathrm{dom}(\mathbf{E})$ then
9:   $\textbf{O}^{k}\leftarrow\textbf{O}^{k}\cup o$
10:   $\textbf{U}\leftarrow\textbf{U}\cup(\mathcal{ID}_{o},\mathcal{ID}_{{o}^{r}})$
11:   $\textbf{I}^{total}\leftarrow\textbf{I}^{total}\cup\mathbf{E}[(\mathcal{ID}_{o},\,\mathcal{ID}_{{o}^{r}})]$
12:  end if
13:end for
14:while $\textbf{U}\neq\varnothing$ do $\triangleright$ Greedy selection
15:  $\mathcal{I}^{\star}\leftarrow\arg\max_{\mathcal{I}\in\textbf{I}^{total}}\ |\mathbf{H}[\mathcal{I}]\cap\mathbf{U}|$
16:  $\mathbf{E}^{k}[\mathbf{H}[\mathcal{I}^{\star}]\cap\mathbf{U}]\leftarrow\mathcal{I}^{\star}$ $\triangleright$ $\mathcal{I}^{\star}$ cover the most edges.
17:  $\textbf{U}\leftarrow\textbf{U}\setminus\mathbf{H}[\mathcal{I}^{\star}]$ $\triangleright$ Remove covered edges
18:end while
[…]
20:return $\textbf{S}^{k}$
[…]
  • Focus. The compressed scene graph $\hat{\textbf{S}}$ significantly reduces the overhead, therefore, we directly fed it to VLM to select the top‑k related objects $\textbf{O}^{rel}\subset\textbf{O}$ where most relevant to the current navigation target.
  • Pruning. Finally, we construct the key subgraph $\textbf{S}^{k}=(\textbf{O}^{k},\textbf{E}^{k})$ by progressively selecting the images that cover the most edges through a dynamic greedy allocation algorithm (Algo.  1).
@@
This technique substantially reduces the information fed to VLM for reasoning, saving token costs by over 95% while retaining most of the important objects relevant to the task. Consequently, the agent can pay more attention to task‑relevant entities more efficiently, improving both decision speed and overall navigation performance.

### v2

@@ 1 Introduction
Previous RL-based embodied navigation methods [rlnav1, rlnav2] suffer from poor generalization and a large sim-to-real gap [sg-nav]. Inspired by rapid advances in pre-trained large model techniques [gpt4o, qwen], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [3dmem, dynavlm, vlngame, tango] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
@@
  • Visual Deficiency. Converting visual observations into text-only graphs discards visual evidence, leading to ambiguity and sensitivity against perception errors.
  • Constrained Vocabulary. Novel categories beyond a preset vocabulary cannot be represented, limiting generalization in 3D scene graph-based methods.
Figure 2: Performance comparisons between our method and other existing ones for embodied navigation. (a) The superiority of our M3DSG over traditional 3D scene graphs. (b) Distance statistics from the goal for the previous method (3D-Mem [3dmem] as an example). (c) Our MSGNav system achieves state-of-the-art performance on both HM3D-OVON and Goat-Bench benchmarks.
Based on these observations, we conclude that visual information is indispensable to the real-world navigation process. Therefore, as illustrated in Fig. 1, we introduce the Multi-modal 3D Scene Graph (M3DSG), which replaces the pure-text relational edges with dynamically assigned images to incorporate visual cues, and facilitates efficient and open-vocabulary inference. Specifically, our M3DSG effectively addresses the three limitations above:
@@
  • Visual Replenishment: M3DSG provides visual contexts to enhance robustness against perception errors.
  • Unconstrained Vocabulary: M3DSG utilizes preserved visual contexts to expand vocabulary, enabling open-vocabulary generalization dynamically.
@@
Additionally, we realize that the last‑mile problem in navigation tasks — knowing the target’s location is not equivalent to determining a suitable navigation pose [^1] Cannot see the wood for the trees. Being too close or occluded prevents a clear observation of the target, which is not a good viewpoint.. Existing zero-shot embodied navigation methods usually stop before that final step. As shown in Fig. 2(b), the existing 3D-Mem method [3dmem] successfully approaches the target but still fails in many navigation tasks due to unacceptable viewpoints (distances ranging from 0.25 m to 1.0 m). To address this issue, we introduce a visibility‑based viewpoint decision module in our MSGNav.
@@
  • M3DSG: We propose a multi-modal 3D scene graph that incorporates visual information, overcoming pure-text limitations and enhancing open-vocabulary scene representation for embodied navigation.
  • Last-Mile: We identify the last-mile problem and propose a visibility‑based viewpoint decision module to address it.
  • MSGNav: We develop a zero-shot embodied navigation system built on M3DSG, enabling active exploration and achieving state-of-the-art performance on the GOAT-Bench and HM3D-OVON embodied navigation datasets.
@@ 3 Approach
Figure 3: The overall framework of our MSGNav. At time step $t$, the agent incrementally constructs the scene graph $S_{t}$ based on received observation $\mathcal{I}_{t}$ and its own pose. $S_{t}$ includes a set of objects $\textbf{O}_{t}$ with attributes, namely visual, spatial, and room properties, along with a set of image edges $\textbf{E}_{t}$ representing object relationships. Subsequently, $S_{t}$ is processed through KSS, AVU, and CLR modules, before being input to VLM query to obtain the target object $\bar{o}$. Finally, VVD module selects the insightful viewpoint $\textbf{v}_{best}$ as a navigation point.
@@ 3.1.1 Problem definition
We consider a mobile embodied agent operating in an unknown or partially explored environment (i.e. the lifelong navigation setting). The agent receives target information in the form of an category, natural language description, or reference image. At each time step $t$, it obtains an RGB-D observation $\mathcal{I}_{t}$ and executes an action $\mathcal{A}_{t}$ (camera rotation or ego-motion) to actively explore until locating the target. The task is successful if the agent reaches any target viewpoint within $d$ meters in at most $T$ steps; otherwise, it fails. In the zero-shot navigation setting, the agent performs navigation without task-specific training or fine-tuning in simulation, thereby reducing training cost and improving generalization to unseen environments and targets.
@@ 3.1.2 Overview
Zero-shot goal-oriented embodied navigation remains highly challenging for conventional learning-based methods. Following current best practices, we utilize off-the-shelf pre-trained models: Vision Foundation Models (VFMs) and Vision Language Models (VLMs) for primitive reasoning. While these models excel in visual understanding, they are limited in 3D spatial reasoning. To address this gap, we propose the M3DSG, an explicit and multimodal 3D scene graph that is incrementally built to encode scene contexts while preserving the visual evidence. Based on M3DSG, the navigation system MSGNav completes goal-oriented navigation with a Visibility-based Viewpoint Decision (VVD) module.
@@ 3.2 M3DSG: Multi‑modal 3D Scene Graph
Traditional 3D scene graphs represent object relations with text-only edges, which are overly abstract and hinder VLMs from accurately understanding scene contexts. Inspired by 3D-Mem [3dmem], which emphasizes the value of raw images for navigation, we retain visual information when constructing an object-centric 3D scene graph. Unlike traditional 3D scene graph [Conceptgraphs] which uses textual relation edges, our method stores images to describe detailed object relations directly. This image edge preserves the benefits of 3D scene graphs while avoiding repeated, costly model queries and providing a more holistic scene representation.
@@ 3.2.1 M3DSG Structure Definition
During the environment exploring process, M3DSG constructs a scene graph $\mathbf{S}=(\mathbf{O},\mathbf{E})$, where $\mathbf{O}=\{o_{i}\}_{i=1}^{N_{o}}$ and $\mathbf{E}=\{\mathbf{e}_{j}\}_{j=1}^{N_{e}}$ denote the sets of objects and edges, respectively. Here, $N_{o}$ is the number of detected objects, and $N_{e}=\sum_{1\leq x<y\leq N_{o}}\mathds{1}_{\|Pos_{o_{x}}-Pos_{o_{y}}\|<\theta}$ is the number of object pairs $(o_{x},o_{y})$ whose spatial distance is below $\theta$. Each object $o_{i}$ is described by attributes including corresponding category, 3D coordinates, and room locations. Each edge $\mathbf{e}_{j}$ stores a set of RGB-D images $\mathbf{I}_{j}=\{\mathcal{I}_{k}\}_{k=1}^{|\mathbf{I}_{j}|}$ that represent the pair of corresponding objects to record their visual relationship.
@@ 3.2.2 Incremental Construction of M3DSG
Following prior work [sg-nav, unigoal], the scene graph $\mathbf{S}$ is incrementally updated during exploration. At step $t$, $\mathbf{S}_{t-1}$ is updated to $\mathbf{S}_{t}$ with the current RGB-D observation $\mathcal{I}_{t}$ as:
@@
Object Update. For the RGB-D observation $\mathcal{I}_{t}$ at time $t$, we extract the frame-level object set $\mathbf{O}^{frame}_{t}=\{o_{i}\}_{i=1}^{|\mathbf{O}^{frame}_{t}|}$ by VFMs: YOLO-W [yolo] for open-vocabulary detection, SAM [sam] for instance masks, and CLIP [clip] for visual embeddings. Each object $o_{i}$ is represented as $o_{i}=\{\mathcal{ID},\mathcal{C},\mathcal{P},\mathcal{B},\mathcal{M},\mathcal{PC},\mathcal{V},\mathcal{R}\}_{i},$ denoting unique ID, category, 3D position, bounding box, semantic mask, point cloud, visual feature, and room location, respectively. The global object set $\mathbf{O}_{t}$ is incrementally updated from the previous set $\mathbf{O}_{t-1}$ and the frame-level object set $\mathbf{O}^{frame}_{t}$ as:
@@
Edge Update. The core of an edge is a set of images $\mathbf{I}$ in which an object pair co-occurs, indexed by the unique IDs of the object pair, i.e. $\mathbf{E}:(\mathcal{ID},\mathcal{ID})\rightarrow\mathbf{I}.$
Meanwhile, we maintain a mapping from each image to its corresponding object pair, i.e. $\mathbf{H}:\mathcal{I}\rightarrow(\mathcal{ID},\mathcal{ID}).$
For the current RGB-D frame $\mathcal{I}_{t}$ at step $t$ with object set $\mathbf{O}^{frame}_{t}$ (from Object Update), we generate co-occurring object pairs $(o_{x},o_{y})$ that appear together in $\mathcal{I}_{t}$ as:
@@
where $\theta$ is the adjacency distance threshold. The edge set $\mathbf{E}_{t}$ is then updated by retrieving the unique IDs $(\mathcal{ID}_{o_{x}},\mathcal{ID}_{o_{y}})$ of each detected pair and appending $\mathcal{I}_{t}$ to the corresponding image set as:
@@
This edge update process is efficient, eliminating the need for costly VLM queries. Moreover, the stored image recorded in $\mathbf{E}_{t}$ provides compact and visually grounded representations of object relationships, which can be directly utilized for downstream exploration reasoning.
@@ 3.3 MSGNav Embodied Navigation System
M3DSG provides comprehensive scene representations that preserve rich visual information. To fully exploit this, we propose the navigation system MSGNav. As shown in Fig.  3(b), our MSGNav first significantly reduces the tokens and time cost required for inference by selecting key subgraphs (Sec. 3.3.1). MSGNav then addresses the out-of-vocabulary issue inherent to pre-set vocabulary by adaptively updating vocabulary with visual evidence from M3DSG (Sec. 3.3.2). Meanwhile, MSGNav performs closed-loop reasoning via decision memory and feedback (Sec. 3.3.3). Finally, MSGNav tackles the “last-mile” problem by optimizing navigation points through visibility-based viewpoint decisions (Sec. 3.3.4).
@@ 3.3.1 Key Subgraph Selection (KSS)
During exploration, the 3D scene graph $\mathbf{S}$ grows progressively and can become voluminous, thereby impairing the decision-making efficiency [3dmem]. In embodied navigation, however, typically only a small related subgraph to the required target is meaningful. Therefore, processing the full scene graph consumes substantial computational resources without commensurate benefits. The key to addressing this issue is to preserve maximal relevant information with the minimal scene graph. MGSNav extracts the target‑relevant subgraphs via a Compress–Focus–Prune procedure:
@@
6:for $o\in\mathbf{O}$ do $\triangleright$ Filter relevant edges
7:  if $\exists~{o}^{r}\in\textbf{O}^{rel}$: $(\mathcal{ID}_{o},\,\mathcal{ID}_{{o}^{r}})\in\mathrm{dom}(\mathbf{E})$ then
8:   $\textbf{O}^{k}\leftarrow\textbf{O}^{k}\cup o$
9:   $\textbf{U}\leftarrow\textbf{U}\cup(\mathcal{ID}_{o},\mathcal{ID}_{{o}^{r}})$
10:   $\textbf{I}^{total}\leftarrow\textbf{I}^{total}\cup\mathbf{E}[(\mathcal{ID}_{o},\,\mathcal{ID}_{{o}^{r}})]$
11:  end if
12:end for
13:while $\textbf{U}\neq\varnothing$ do $\triangleright$ Greedy selection
14:  $\mathcal{I}^{\star}\leftarrow\arg\max_{\mathcal{I}\in\textbf{I}^{total}}\ |\mathbf{H}[\mathcal{I}]\cap\mathbf{U}|$
15:  $\mathbf{E}^{k}[\mathbf{H}[\mathcal{I}^{\star}]\cap\mathbf{U}]\leftarrow\mathcal{I}^{\star}$ $\triangleright$ $\mathcal{I}^{\star}$ cover the most edges.
16:  $\textbf{U}\leftarrow\textbf{U}\setminus\mathbf{H}[\mathcal{I}^{\star}]$ $\triangleright$ Remove covered edges
17:end while
18:return $\textbf{S}^{k}=(\textbf{O}^{k},\textbf{E}^{k})$ $\triangleright$ Key scene graph
@@
  • Focus. The compressed scene graph $\hat{\textbf{S}}$ significantly reduces the overhead. Therefore, we directly fed it to the VLM to select top‑k related objects $\textbf{O}^{rel}\subset\textbf{O}$ which are the most relevant to the current navigation target.
  • Pruning. Finally, we construct the key subgraph $\textbf{S}^{k}=(\textbf{O}^{k},\textbf{E}^{k})$ by progressively selecting the image that covers the most edges through a dynamic greedy allocation algorithm (in Algorithm  1).
@@
This technique reduces the token cost by over 95% while feeding the information to VLM for reasoning, retaining most of the important objects relevant to the task. Consequently, the agent can efficiently focus on task‑relevant entities, improving both navigation speed and performance.